### v2

@@ Abstract
Two office-driven politicians compete in an election by proposing policies. There are two possible states of the world: climate change is either mild, with no lasting effect on welfare if addressed properly, or severe, leading to reduced welfare even with appropriate measures. Voters receive signals about the state but may interpret them in a non-Bayesian way, holding motivated beliefs. An equilibrium always exists where voters ignore signals suggesting severe consequences, causing politicians to propose policies for mild climate change—even when they know otherwise. If severe climate change leads to only moderate welfare losses, an efficient equilibrium also exists. In this equilibrium, voters trust politicians to choose the optimal policies, implying voters choose to trust their signals, which in turn encourages optimal policy choices by politicians. The model highlights the role of political rhetoric and trust in government, and a first glance at the data reveals patterns consistent with the models predictions.
@@
Keywords: political competition, climate change, motivated reasoning, trust in government, political rhetoric, misinformation
@@ 1 Introduction
Clayton, Manning, Krygsman, and Speiser (2017)
“Yes, There Has Been Progress on Climate.
No, It’s Not Nearly Enough.”
The New York Times, 25 Oct 2021
@@
Climate change is one of the most pressing issues of our time, threatening the livelihoods of millions of people around the globe. Information about climate change and global warming has been available for more than a century, dating back at least to Svante Arrhenius’ (1896) famous paper “On the Influence of Carbonic Acid in the Air upon the Temperature of the Ground.” However, despite an ever-growing scientific consensus that increasing atmospheric CO2 concentrations will severely impact our planet, little to no action was taken to stop this process for most of the last few decades. Many reasons have been put forward to explain the inaction of political decision-makers, ranging from lobbying interests to widespread misinformation campaigns like those of ExxonMobil. In this paper, I discuss a different, simpler channel: the electoral incentives of politicians when voters may choose to hold motivated beliefs.
@@
The looming dire consequences of climate change may, if taken seriously, create stress and anxiety. In an effort to avoid these negative emotions, people may choose to hold motivated beliefs, ignoring information that suggests climate change is severe, while overreacting to information suggesting there is nothing to worry about. Psychologists and economists alike have long been aware of such information processing biases, as exemplified by Kunda’s (1990) seminal paper or the recent survey by Amelio and Zimmermann (2023), and multiple studies have provided evidence for their empirical relevance (for example, Lewandowsky, Gignac, and Oberauer, 2013, Thaler, 2021, 2024, and Lois, Tsakas, Yuen, and Riedl, 2024). To study the potential consequences of such behavior for the implementation of policies targeted at climate change, I construct a game-theoretic model in which two politicians compete in an election by proposing policies. There are two possible states of the world: severe climate change and moderate (or no) climate change. The state determines which policy is optimal and establishes a baseline welfare level. If climate change is severe, the baseline welfare is lower than in the mild state, regardless of the chosen policy. Voters and politicians receive signals about the state of the world, and society is best served when the policy that matches the state is implemented. As discussed by authors like Bénabou and Tirole (2016) or Spiegler (2016), voters have anticipatory utility, meaning they feel anxiety or stress when the future looks grim. To cope with these negative emotions, they may choose to hold motivated beliefs, interpreting information about the future in a non-Bayesian way to increase anticipatory utility. How voters react to information is an important determinant of how politicians compete to win an election.
The model highlights the importance of the portrayed severity of climate change for the efficiency of policy-making. In particular, if severe climate change has only moderate consequences for baseline welfare levels, then there always exists an efficient equilibrium in which politicians choose the optimal policies for each state. However, this equilibrium only exists if voters have sufficient trust in government, i.e., if they belief that elected politicians will enact the right policies with a sufficiently high probability. Otherwise, with low trust in government, there always exists an equilibrium in which politicians ignore their information and choose policies for mild or no climate change.
The analysis suggests two important determinants for the efficiency of policy-making. First, trust in government is important, because it determines in which way voters interpret information when severe climate change only has moderate effects on baseline welfare. With sufficient trust in government, there always exists an efficient equilibrium in which politicians strictly choose the optimal policies. To the contrary, if trust in government is low, then politicians will ignore their information and policy choice will be inefficient. Second, the analysis suggests that political rhetoric could be an important determinant of the quality of policy choice as well. In particular, if the consequences of severe climate change are portrayed as catastrophical, then the efficient equilibrium ceases to exist. Voters will ignore any information that climate change is severe, inducing office motivated politicians to ignore their information as well, and in the unique equilibrium politicians always choose policies for mild or no climate change. However, if the consequences of severe climate change are portrayed as moderate or mild, then the efficient equilibrium survives.
Literature:
@@
The paper contributes to several areas of literature. Firstly, it adds to the literature that examines the incentives of political candidates to select efficient policy platforms when the true state of the world is unknown. Important contributions to this literature include Heidhues and Lagerlof (2003), Laslier and van der Straeten (2004), and Kartik, Squintani, and Tinn (2015). Unlike Heidhues and Lagerlof (2003) and Kartik, Squintani, and Tinn (2015), I assume that voters also receive information, which creates the main tension of whether to ignore it or not. In Laslier and van der Straeten (2004), voters receive a single public signal as well, whereas in the current paper, each voter receives an independent signal. The two most closely related papers in this literature are Gratton (2014) and Crutzen, Sisak, and Swank (2024). Gratton (2014) assumes that candidates observe the state perfectly while voters receive imperfect but informative signals about the state, and the focus of the analysis is to identify conditions that lead equilibria such that candidates choose efficient policies. Crutzen, Sisak, and Swank (2024) study a game like Gratton (2014), but there are two groups of voters, elites and commoners, who stochastically differ in their policy preferences and who always differ in quality of the information they receive. The analysis focusses on deriving conditions such that candidates cater to the elites or engage in populism. Unlike in those papers, the focus in the current paper is on misinterpreting information, and how this influences candidates’ incentives to provide efficient policies.
@@
A cornerstone of the model is that voters have anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare” (Bénabou and Tirole, 2016). As in Akerlof and Dickens (1982), Caplin and Leahy (2001), Bénabou and Tirole (2002), Spiegler (2016), Eliaz and Spiegler (2020), and other papers, this means that voters have preferences not only about states and policies but also about their own beliefs. A consequence of this is that voters may engage in motivated reasoning; see, for example, Kunda (1990), Bénabou and Tirole (2016), Zimmermann (2020), Oprea and Yuksel (2021), or Thaler (2024). The most closely related papers in this area are Levy (2014) and Le Yaouanq (2023), which also examine electoral contexts. Levy (2014) presents a model featuring a policymaker trying to signal quality/congruency to voters. Voters have imperfect memory and can suppress certain news by purposefully conflating positive and negative signals. In contrast, there is no signaling in the current paper. Furthermore, voters have perfect memory but may choose to interpret some signals as more or less informative than they really are. Finally, it is not just voters’ beliefs about the state that matter, but also their expectations regarding policy outcomes. This difference introduces a novel self-fulfilling aspect that becomes critical for equilibrium. Le Yaouanq (2023) suggests that voters may interpret signals in a contrary manner at a cost. In the current paper, I assume that voters can modulate the perceived informativeness of a signal, but they cannot alter its direction. Moreover, unlike in Le Yaouanq (2023), a key component of my model is that the state influences baseline welfare levels, which has important consequences for equilibrium.
@@
The paper also contributes to a recently revived literature on environmental policymaking. Delfgaauw and Swank (2024) show that there can be locked-in effects that prevent the adoption of environmentally friendly policies. Blumenthal (2024) shows that efficient environmental policies may not be adopted if voters’ preferences are expected to change over time. Van der Straten, Perotti, and Van der Ploeg (2024) study the relationship between climate change adoption policies and inequality in society. Besley and Persson (2023) examine the conditions for green transitions to occur, depending on voters’ ideological preferences and the extent to which the production sector remains reliant on brown energy sources. Gullberg (2008), Shapiro (2016), and Balles, Matter, and Stutzer (2024) show that special interest groups may prevent climate change policies from being adopted. Denter (2024) shows that the wrong kind of climate activism may decrease support for climate policies. In contrast to these papers, the current paper provides an explanation based purely on the electoral incentives of office-motivated politicians in the presence of a common information processing bias among voters.
@@
The paper is organized as follows. In Section 2, I describe the baseline model, which is then solved in Section 3. Section 4 extends the baseline model along several dimensions. In Section 5, I derive testable predictions of the model and have a first quick look at the data. Section 6 concludes. All proofs are relegated to the appendix.
@@ 2 The Model
In this section I introduce the main building blocks of the model. There is an unknown state of the world, $\omega\in\{0,1\}$, indicating whether climate change is mild, $\omega=0$, or severe, $\omega=1$. The prior probability that $\omega=1$ is $q\in(0,1)$. There are two different kinds of actors. On the one hand, there are two kinds of purely office motivated politicians, indexed by $i=1,2$. On the other hand, there is a continuum of voters of mass 1. Candidates vie for voters’ support by proposing policies, $p_{i}\in\{0,1\}$, to which they commit. Denote the vector of policies by $\mathbf{p}\equiv(p_{1},p_{2})$.
@@
$u(p,\omega)=\left\{\begin{array}[]{cl}-p&\text{if }\omega=0,\\ -\Delta-\beta\left|p-1\right|&\text{if }\omega=1.\end{array}\right.$ (1)
Hence, for each state realization $\omega$, the optimal policy is $p=\omega$. But the state $\omega$ determines not only optimal policy, but the baseline welfare level in society. In particular, if $\omega=1$, i.e., when climate change is severe, then baseline welfare decreases by $\Delta>0$, independent of the chosen policy, while it remains constant if climate change is mild, $\omega=0$. The parameter $\beta>0$ measures the importance of choosing the right policy when climate change is severe: the greater is $\beta$, the more important is taking the right action. Note that the worst outcome is choosing policy incorrectly in state 1, yielding $-\Delta-\beta$, while the best outcome is choosing correctly in state 0, yielding a utility of zero. I make the following assumption to capture the idea that at the outset, voters believe the optimal policy choice is $p=0$: [^1] This assumption is not without loss of generality for the model’s conclusions. Allowing for $q\geq\frac{1}{1+\beta}$ would imply that the set of possible equilibria is larger. I chose to assume $q<\frac{1}{1+\beta}$, because it seems realistic that voters need to be convinced to support anti-climate change policies. Moreover, as a second benefit, Assumption 1 keeps the paper’s exposition more concise.
@@
Throughout I assume that $q<\frac{1}{1+\beta}.$
At the beginning of the game, both types of players receive a signal that is informative about $\omega$. In particular, as in Gratton (2014), before announcing policy platforms every politician $i$ receives a signal $s_{i}^{P}$ that perfectly reveals $\omega$ Hence, after receiving $s_{i}^{P}$, each politician perfectly knows wether climate change is severe or not.
Voters receive two kinds of signals. On the one hand, the policy platforms $\mathbf{p}$ may function as signals about $\omega$, depending on the strategies chosen by the candidates. On the other hand, they receive information from news or from direct experience about $\omega$. I model this by assuming that each voter receives a signal $s\in\mathbb{R}$. When the state is $\omega$, the c.d.f. of $s$ is $\Phi\left(\frac{s+(2\omega-1)\mu}{\sigma}\right)$, where $\Phi$ is the c.d.f. of the standard Gaussian distribution and $\mu>0$ and $\sigma>0$ are parameters. Therefore, the typical monotone likelihood ratio property (MLRP) is satisfied, meaning that any $s>0$ is evidence for $\omega=1$ and increases the belief compared to the prior, while any $s<0$ is evidence for $\omega=0$ and decreases the belief. Following Callander (2011) and others, I interpret $\mu$ as a measure of the complexity of the issue climate change. If $\mu$ is small, the issue is very complex, and thus voters tend to hold imprecise beliefs about $\omega$. To the contrary, if $\mu$ is large, climate change is not a very complex issue and beliefs tend to be precise.
@@
Throughout I assume that
$\mu\geq\mu^{\text{Bayes}}\equiv\frac{\sigma\sqrt{\ln\left(\frac{1-q}{\beta q}\right)}}{\sqrt{2}}.$
I assume that voters have anticipatory utility. In particular, a voter has expectations about her future utility, which is derived from a belief bout the true state and a second belief about the enacted policy. This causes anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare(Bénabou and Tirole, 2016). To increases anticipatory utility (to be defined precisely below), a voters may use motivated reasoning, which means that she may update beliefs using a distorted complexity parameter $\tilde{\mu}$. Clearly, processing information using such a distorted complexity parameter comes at a cost, because higher anticipatory utility may imply lower utility because of imperfect decision making. As Bénabou and Tirole (2016) write, “one can react to bad news objectively, which leads to better decisions but having to live with grim prospects for some time […], or adopt a more “defensive” cognitive response that makes life easier until the day of reckoning, when mistakes will have to be paid for.” I do not model these costs in the baseline model to be able to identify the pure effect of motivated reasoning, but in Section 4.4 two models of distortions are discussed. If a voter is indifferent between having $\tilde{\mu}=\mu$ and some other $\tilde{\mu}\neq\mu$, I assume without loss of generality that she chooses $\tilde{\mu}^{*}=\mu$. Choosing $\tilde{\mu}>\mu$ implies the voter interprets the signal as more informative than it really is, and hence beliefs will change excessively, while choosing $\tilde{\mu}<\mu$ implies a more conservative stance and that beliefs move less than they should.
Based on $s$ and $\mathbf{p}$, and after choosing $\tilde{\mu}$, each voter forms beliefs $\pi(s,\mathbf{p},\tilde{\mu})$ and $\kappa$. $\pi(s,\mathbf{p},\tilde{\mu})$ is the posterior belief about $\omega$, whereas $\kappa$ is the probability assigned to the event that policy 1 will be implemented after the election. Because candidates commit to policy platforms, clearly $\kappa=1$ if $p_{1}=p_{2}=1$, and $\kappa=0$ if $p_{1}=p_{2}=0$.
@@
$AU(s,\mathbf{p},\tilde{\mu})=-\kappa\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot\Delta+(1-\pi(s,\mathbf{p},\tilde{\mu}))\right]-(1-\kappa)\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot(\Delta+\beta)\right]$ (2)
@@
Note that a voter is never pivotal, because there is a continuum of voters. I thus assume that voters vote sincerely as defined by Austen-Smith and Banks (1996). In words, each voter forms beliefs based on $s$, $\mathbf{p}$, and $\tilde{\mu}$ and then votes to maximize (1).
We are now in a position to define equilibrium of our game. (i) each voter maximizes $AU(s,\mathbf{p},\tilde{\mu})$ given a belief about the probability that policy 1 will be enacted $\kappa$; (ii) each voter votes sincerely given $s$ and $\tilde{\mu}$; (iii) $\kappa$ is correct given $s$, $\tilde{\mu}$, and given the strategies of all other voters; (iv) politicians choose policy platforms that maximize the probability to be elected, given their expectations about voters’ behavior. As typical in signaling games, there are multiple equilibria. I focus attention on symmetric equilibria, that is, equilibria in which candidates, holding the same information and trying to achieve the same, choose identical strategies. Further, if on the equilibrium path both candidates choose the same pure strategy, then we must have $p_{1}=p_{2}$. If, off the equilibrium path, voters observe $p_{1}\neq p_{2}$, I assume that they do not learn anything from this about $\omega$. That is, if off the equilibrium path $p_{1}\neq p_{2}$, then observing $\mathbf{p}$ does not change voters’ belief about $\omega$ [^2] This could be justified, for example, because candidates have types as well. An ideological candidate always chooses $p_{i}=i-1$, $i\in\{0,1\}$. A strategic candidate chooses the policy that is expected to maximize the probability to win the election. A model in which each candidate $i$ is ideological with probability $\gamma\in(0,1)$ and non-ideological with probability $1-\gamma$ would yield the same results regarding the platform choices of office-motivated candidates as the assumption of non-informative deviations.

### v5

@@ Abstract
We study how motivated reasoning affects the provision of climate policy in an electoral competition framework. Voters experience anticipatory disutility when future outcomes appear grim and may therefore distort beliefs in response to adverse information. We develop a game-theoretic model in which voters and politicians receive signals about the severity of climate change. When the anticipated welfare losses from severe climate change are sufficiently large, voters optimally ignore unfavorable information, inducing politicians to campaign on policies appropriate for mild climate change only. When welfare losses are moderate, the model admits a second, efficient equilibrium in which voters trust politicians to implement welfare-maximizing policies and vote informatively, thereby creating incentives for politicians to propose adequate climate policy. The model shows how motivated belief formation and voters’ expectations about policy responsiveness jointly determine equilibrium selection between effective climate policy and persistent political inaction.
@@
Keywords: political competition, climate change, motivated reasoning, trust in government, political rhetoric
@@ 1 Introduction
Clayton et al., (2017)
@@
Climate change is one of the most pressing issues of our time, threatening the livelihoods of millions of people around the globe. Information about climate change and global warming has been available for more than a century, dating back at least to Svante Arrhenius’ (1896) famous paper “On the Influence of Carbonic Acid in the Air upon the Temperature of the Ground.” However, despite an ever-growing scientific consensus that increasing atmospheric CO2 concentrations will severely impact our planet, little to no action was taken to stop this process for most of the last few decades. Many reasons have been put forward to explain the inaction of political decision-makers, ranging from lobbying interests to widespread misinformation campaigns like those of ExxonMobil (e.g., Oreskes and Conway,, 2011, or Supran et al.,, 2023). In this paper, we discuss a different, simpler channel: the electoral incentives of politicians when voters may hold motivated beliefs.
@@
The looming dire consequences of climate change may, if taken seriously, create stress and anxiety. In an effort to avoid these negative emotions, people may choose to hold motivated beliefs, ignoring information that suggests climate change is severe, while overreacting to information suggesting there is nothing to worry about. Psychologists and economists alike have long been aware of such information processing biases, as exemplified by Kunda’s (1990) seminal paper or the recent survey by Amelio and Zimmermann, (2023), and multiple studies have provided evidence for their empirical relevance (for example, Lewandowsky et al.,, 2013, Thaler,, 2021, 2024, and Lois et al.,, 2024).
Motivated reasoning alone, however, is unlikely to fully account for persistent political inaction on climate change. Recent empirical work has also emphasized the role of political trust in shaping climate policy. In particular, higher trust in political institutions has been associated with local politicians adopting more ambitious climate policies (Bose,, 2023). Figure 1 shows that a similar positive relationship between political trust and climate policy stringency emerges in a cross-country comparison. [^1] Data on climate policy performance come from the Climate Change Performance Index (CCPI), published jointly by Germanwatch, the NewClimate Institute, and the Climate Action Network International. The CCPI ranges from 0 to 100, with higher values indicating better performance in climate policy. It aggregates indicators on greenhouse gas emissions (40%), renewable energy (20%), energy use (20%), and other climate policies (20%). Data on trust are taken from the OECD’s Trust in Government Indicator (TGI), which measures the share of a country’s population reporting confidence in their national government and also ranges from 0 to 100. These patterns point to an interaction between belief distortions and political trust in determining policy outcomes, a mechanism that the analysis below formalizes.
Figure 1: Scatterplot with fitted trendline and 95% confidence interval showing countries’ CCPI scores on the vertical axis and TGI scores on the horizontal axis.
The aim of this paper is to construct a game-theoretic model that sheds light on how motivated reasoning affects politicians’ incentives to propose adequate climate change policies within an electoral competition framework. In the model, there are two possible states of the world: severe climate change and mild climate change. The state determines which policy is optimal and establishes a baseline level of welfare. If climate change is severe, baseline welfare is lower than in the mild state, regardless of the policy chosen. Voters and politicians receive signals about the state of the world, and society is best served when the implemented policy matches the true state. As in Brunnermeier and Parker, (2005), Bénabou and Tirole, (2016), or Spiegler, (2016), voters derive anticipatory utility and experience anxiety or stress when the future appears grim. To cope with these negative emotions, they may adopt motivated beliefs, interpreting information about the future in a non-Bayesian manner to thus increase anticipatory utility. How voters respond to information is therefore a key determinant of how politicians compete in elections.
The analysis reveals that when the consequences of severe climate change are sufficiently large (“catastrophic”), voters maximize anticipatory utility by disregarding signals indicating that the problem is severe. This, in turn, induces politicians to campaign exclusively on policies designed for mild climate change. In contrast, when the welfare losses from severe climate change are moderate rather than catastrophic, multiple equilibria arise. In an efficient equilibrium, voters trust politicians to choose welfare-maximizing policies, thereby creating incentives for politicians to do so. In an inefficient equilibrium, voters expect politicians to ignore relevant information and to campaign on policies appropriate for mild climate change, which in turn gives politicians incentives to disregard severe climate risks. Thus, when the consequences of severe climate change are moderate, voter trust in welfare-maximizing policy choices becomes self-fulfilling, sustaining politicians’ incentives to propose such policies.
Trust in government functions as a coordination device. When voters believe that politicians will implement policies that match the underlying state, signals are interpreted more faithfully, allowing society to converge to the efficient equilibrium. By contrast, when trust is low—interpreted as the belief that policy choices are largely independent of the underlying state—this link breaks down. In that case, pessimistic voters interpret information in ways that rationalize the policies they expect the government to pursue, giving rise to a self-confirming equilibrium in which inefficient policies persist. The model therefore identifies trust as an important determinant of whether political institutions can translate information about climate risks into effective action.
The analysis shows that an efficient equilibrium is less likely to exist when the anticipated welfare losses from severe climate change are sufficiently large. How the severity of climate change is framed—whether it is presented as a “catastrophe” or as a more moderate “challenge”—may influence voters’ expectations about the magnitude of potential losses from climate change. An indirect implication of the model is, therefore, that climate change rhetoric can influence voters’ anticipatory utility and, through this channel, their belief formation, which in turn determines equilibrium outcomes. In this sense, highly alarming descriptions need not translate into greater support for policy action and may instead induce defensive belief formation, a pattern emphasized by Clayton et al., (2017) in the quote at the beginning of the paper.
Related Literature:
@@
The paper contributes to several areas of literature. Firstly, it adds to the literature that examines the incentives of political candidates to select efficient policy platforms when the true state of the world is unknown. Important contributions to this literature include Heidhues and Lagerlof, (2003), Laslier and van der Straeten, (2004), and Kartik et al., (2024). Unlike Heidhues and Lagerlof, (2003) and Kartik et al., (2024), we assume that voters also receive information, which creates the main tension of whether to ignore it or not. In Laslier and van der Straeten, (2004), voters receive a single public signal as well, whereas in the current paper, each voter receives an independent signal. The two most closely related papers in this literature are Gratton, (2014) and Crutzen et al., (2024). Gratton, (2014) assumes that candidates observe the state perfectly while voters receive imperfect but informative signals about the state, and the focus of the analysis is to identify conditions that lead to equilibria such that candidates choose efficient policies. Crutzen et al., (2024) study a game like Gratton, (2014), but there are two groups of voters, elites and commoners, who stochastically differ in their policy preferences and who always differ in quality of the information they receive. The analysis focusses on deriving conditions such that candidates cater to the elites or engage in populism. Unlike in those papers, the focus of the current paper is on voters’ misinterpretation of information and how this influences candidates’ incentives to propose efficient policies. This is related to Millner et al., (2020), who study how confirmation bias affects electoral outcomes. In contrast, we consider a model in which the future affects voters’ well-being through anticipatory utility, giving them an incentive to hold motivated beliefs.
@@
As just foreshadowed, a cornerstone of the model is that voters derive anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare” (Bénabou and Tirole,, 2016). As in Akerlof and Dickens, (1982), Caplin and Leahy, (2001), Bénabou and Tirole, (2002), Brunnermeier and Parker, (2005), Spiegler, (2016), Eliaz and Spiegler, (2020), Little et al., (2022), and Engelmann et al., (2024), this means that voters have preferences not only about states and policies but also about their own beliefs. A consequence of this is that voters may engage in motivated reasoning; see, for example, Kunda, (1990), Bénabou and Tirole, (2016), Zimmermann, (2020), Oprea and Yuksel, (2021), or Thaler, (2024). The most closely related papers in this area are Levy, (2014) and Le Yaouanq, (2023), which also examine electoral contexts. Levy, (2014) presents a model featuring a policymaker trying to signal quality/congruency to voters. Voters have imperfect memory and can suppress certain news by purposefully conflating positive and negative signals. In contrast, there is no signaling in the current paper. Furthermore, voters have perfect memory but may choose to interpret some signals as more or less informative than they really are. Finally, it is not just voters’ beliefs about the state that matter, but also their expectations regarding policy outcomes. This difference introduces a novel self-fulfilling aspect that becomes critical for equilibrium. Le Yaouanq, (2023) suggests that voters may interpret signals in a contrary manner at a cost. In the current paper, we assume that voters can modulate the perceived informativeness of a signal, but they cannot alter its direction. Moreover, unlike in Le Yaouanq, (2023), a key component of our model is that the state influences baseline welfare levels, which has important consequences for equilibrium.
@@
The paper also contributes to the literature on environmental policymaking. Delfgaauw and Swank, (2024) show that there can be locked-in effects that prevent the adoption of environmentally friendly policies. Blumenthal, (2024) shows that efficient environmental policies may not be adopted if voters’ preferences are expected to change over time. Van der Straten et al., (2024) study the relationship between climate change adoption policies and inequality in society. Besley and Persson, (2023) examine the conditions for green transitions to occur, depending on voters’ ideological preferences and the extent to which the production sector remains reliant on brown energy sources. Gullberg, (2008), Shapiro, (2016), and Balles et al., (2024) show that special interest groups may prevent climate change policies from being adopted. In contrast to these papers, the current paper provides an explanation based purely on the electoral incentives of office-motivated politicians in the presence of a common information processing bias among voters.
Finally, the paper contributes to the literature studying trust as a determinant of policy making and welfare. Dasgupta, (2000) argues that higher levels of trust in society reduce transaction costs and thereby increase welfare. Relatedly, Zak and Knack, (2001) show that higher trust stimulates investment and economic growth. Aghion et al., (2010) demonstrate that lower trust in society causes voters to demand more regulation, and that excessive regulation, in turn, hinders the formation of trust. This mechanism is similar to the current model, where low trust in government leads voters to demand policies not targeted at climate change. However, their model predicts a negative relationship between trust and regulation, whereas in the present paper the relationship between trust and climate policy is positive. This is in line with Bose, (2023), who shows that higher trust induces politicians to provide more public goods, in particular, more climate change policies. Besley and Dray, (2024) show that state effectiveness is greater when citizens trust their government, because higher trust increases voluntary compliance and thereby reduces implementation costs. Ehrmann, (2025) reviews the literature on trust in central banks and conclude that a high level of trust is an important prerequisite for a successful monetary policy.
@@
The paper is organized as follows. In Section 2, we describe the baseline model, which is then solved in Section 3. In Section 4, we discuss the main implication of the model. Section 5 extends the baseline model along several dimensions. Section 6 concludes. All proofs are relegated to the appendix.
@@ 2 The Model
In this section we introduce the main building blocks of the model. There is an unknown state of the world, $\omega\in\{0,1\}$, indicating whether climate change is mild, $\omega=0$, or severe, $\omega=1$. The prior probability that $\omega=1$ is $q\in(0,1)$. There are two different kinds of actors. On the one hand, there are two purely office motivated politicians, indexed by $i=1,2$. On the other hand, there is a continuum of voters of mass 1. Candidates vie for voters’ support by proposing policies, $p_{i}\in\{0,1\}$, to which they commit. Denote the vector of policies by $\mathbf{p}\equiv(p_{1},p_{2})$.
@@
$u(p,\omega)=\left\{\begin{array}[]{cl}-p&\text{if }\omega=0,\\ -\Delta-b\left|p-1\right|&\text{if }\omega=1.\end{array}\right$. (1)
Hence, for each state realization $\omega$, the optimal policy is $p=\omega$. But the state $\omega$ determines not only optimal policy, but the baseline welfare level in society. In particular, if $\omega=1$, i.e., when climate change is severe, then baseline welfare decreases by $\Delta>0$, independent of the chosen policy, while it remains constant if climate change is mild, $\omega=0$. The parameter $b>0$ measures the importance of choosing the right policy when climate change is severe: the greater is $b$, the more important is taking the right action. We make the following assumption to capture the idea that at the outset, voters believe the optimal policy choice is p=0p=0: [^2] This assumption is not without loss of generality for the model’s conclusions. Allowing for $q\geq\frac{1}{1+b}$ would imply that the set of possible equilibria is larger. We chose to assume $q<\frac{1}{1+b}$, because it seems realistic that voters need to be convinced to support anti-climate change policies. Moreover, as a second benefit, Assumption 1 keeps the paper’s exposition more concise.
@@
Throughout we assume that $q<\frac{1}{1+b}.$
At the beginning of the game, both types of players receive a signal that is informative about $\omega$. In particular, as in Gratton, (2014), before announcing policy platforms every politician $i$ receives a signal $s_{i}^{P}$ that perfectly reveals $\omega$. Hence, after receiving $s_{i}^{P}$, each politician perfectly knows whether climate change is severe or not.
Voters receive two kinds of signals. On the one hand, the policy platforms $\mathbf{p}$ may function as signals about $\omega$, depending on the strategies chosen by the candidates. On the other hand, they receive information from news or from direct experience about $\omega$. We model this by assuming that each voter receives a signal $s\in\mathbb{R}$. When the state is $\omega$, the c.d.f. of $s$ is $\Phi\left(\frac{s-(2\omega-1)\mu}{\sigma}\right)$, where $\Phi$ is the c.d.f. of the standard Gaussian distribution and $\mu>0$ and $\sigma>0$ are parameters. Therefore, the typical monotone likelihood ratio property (MLRP) is satisfied, meaning that any $s>0$ is evidence for $\omega=1$ and increases the belief compared to the prior, while any $s<0$ is evidence for $\omega=0$ and decreases the belief. Following Callander, (2011) and others, we interpret $\mu$ as a measure of the complexity of the issue climate change. If $\mu$ is small, the issue is very complex, and thus voters tend to hold imprecise beliefs about $\omega$. To the contrary, if $\mu$ is large, climate change is not a very complex issue and beliefs tend to be precise.
@@
Throughout we assume that
$\mu\geq\hat{\mu}\equiv\frac{\sigma\sqrt{\ln\left(\frac{1-q}{bq}\right)}}{\sqrt{2}}.$
We assume that voters derive anticipatory utility. In particular, a voter has expectations about her future utility, which is derived from a belief bout the true state and a second belief about the enacted policy. This causes anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare(Bénabou and Tirole,, 2016). To increases anticipatory utility (to be defined precisely below), a voters may use motivated reasoning, which means that she may update beliefs using a distorted complexity parameter $\tilde{\mu}$. Clearly, processing information using such a distorted complexity parameter comes at a cost, because higher anticipatory utility may imply lower utility because of imperfect decision making. As Bénabou and Tirole, (2016) write, “one can react to bad news objectively, which leads to better decisions but having to live with grim prospects for some time […], or adopt a more “defensive” cognitive response that makes life easier until the day of reckoning, when mistakes will have to be paid for.” We do not model these costs in the baseline model to be able to identify the pure effect of motivated reasoning, but in Section 5.1 two models of distortions are discussed. If a voter is indifferent between having $\tilde{\mu}=\mu$ and some other $\tilde{\mu}\neq\mu$, we assume without loss of generality that she chooses $\tilde{\mu}^{*}=\mu$. Choosing $\tilde{\mu}>\mu$ implies the voter interprets the signal as more informative than it really is, and hence beliefs will change excessively, while choosing $\tilde{\mu}<\mu$ implies a more conservative stance and that beliefs move less than they should.
[…]
@@
$AU(s,\mathbf{p},\tilde{\mu})=-\kappa\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot\Delta+(1-\pi(s,\mathbf{p},\tilde{\mu}))\right]-(1-\kappa)\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot(\Delta+b)\right]$ (2)
@@
Note that a voter is never pivotal, because there is a continuum of voters. We thus assume that voters vote sincerely. In words, each voter forms beliefs based on $s$, $\mathbf{p}$, and $\tilde{\mu}$ and then votes to maximize (1).
[…]